OpenStudy (anonymous):

How can you solve the equation y minus two point nine equals five point three?

OpenStudy (anonymous):

y-2.5 = 5.3 +2.5 = +2.5 ------------ y = 7.8

OpenStudy (anonymous):

So add 2.9 to both sides?

OpenStudy (anonymous):

whoopsie typo, make the 2.5 into 2.9 so y = 8.2